# Conformal Prediction with Corrupted Labels: Uncertain Imputation and Robust Re-weighting

### v2

@@ A.3 Triply robust calibration
Under the assumptions of Theorem 1 we get $\mathbb{P}(Y^{\text{test}}\in C^{\texttt{PCP}}(X^{\text{test}}))\geq 1-\alpha$, under the assumptions of Theorem 4 we get $\mathbb{P}(Y^{\text{test}}\in C^{\texttt{PCP}}(X^{\text{test}}))\geq 1-\alpha$. By [1, 28], CP constructs prediction sets with the nominal coverage level when the scores are exchangeable, i.e., $\mathbb{P}(Y^{\text{test}}\in C^{\texttt{CP}}(X^{\text{test}}))\geq 1-\alpha$. Thus, when at least one of the above assumptions hold, one of the above prediction sets achieves the desired coverage rate. Since TriplyRobust is the union of all three sets, its coverage rate is greater than or equal to the coverage rate of each prediction set. Therefore:
@@ Appendix B Additional related work
Several recent works extend conformal prediction beyond the i.i.d. assumption to handle various forms of distribution shifts, including arbitrary shifts [15], online and time-series settings [29, 30, 31, 32, 33], missing covariates [34, 35], and ambiguous labels [36, 37]. Related research on missing-data imputation includes analyses of missingness mechanisms [38], and multiple imputation frameworks [39, 40]. Furthermore, the work of [41] studies regression imputation approaches and discusses adding residual variance to account for prediction uncertainty.
@@ C.1 Privileged conformal prediction
Compute the scores $S_{i}=\mathcal{S}({X}_{i},\tilde{Y}_{i};\hat{f})$ for the calibration samples, $i\in\mathcal{I}_{2}$.
@@ C.3 Uncertain imputation
Compute the errors $E_{i}$ of $\hat{g}$ on the reference set $\mathcal{I}_{3}$ according to (11).
Generate the imputed labels $\bar{Y}_{i},\forall i\in\mathcal{I}_{2}$, according to (12).
Compute the scores $\bar{S}_{i}=\mathcal{S}({X}_{i},\bar{Y}_{i};\hat{f})$ using the imputed labels for the calibration samples, $i\in\mathcal{I}_{2}$.
@@ C.3.1 Error sampling techniques
In this section, we present our suggestions for sampling errors conditional on the privileged information for the imputation process of UI in (12). Instead of using a strict equality condition of $Z_{i}=z$ in $\mathcal{E}(z):=\{E_{i}:i\in\mathcal{I}_{3},Z_{i}=z\}$, we can relax this requirement by clustering the $Z$ space and comparing the clusters into which each $z$ falls, that is, $\mathcal{E}(z):=\{E_{i}:i\in\mathcal{I}_{3},h(Z_{i})=h(z)\}$, where $h$ is a clustering function. The first clustering method we use is Kmeans, which we fit on the training and validation data. Then, we cluster the PIs of the reference set. When imputing the corrupted labels of the calibration set, we sample an error from the cluster corresponding to the test PI. The second clustering approach we examine is Linear clustering. We first fit on the training and validation data a linear model that takes as an input the PI $Z$ and outputs an estimated label $Y$. We used this model to compute the estimated labels for each point in the reference set and then split them into bins. We note that the alternative approach of learning the error distribution given the PI $Z$ may be applied as well, such as random forest for conditional density estimation [42], or normalizing flows [43, 44].
@@ C.4 Method validity conditions summary
Quantile Regression [45, 46, 47, 48] $\checkmark$ NA NA
@@ D.1 General real dataset details
Dataset # Samples $\boldsymbol{X/Z/Y}$ Dimensions $\boldsymbol{Z}$ description
facebook1 [22] 40948 52/1/1 Number of posts comments
facebook2 [22] 81311 52/1/1 Number of posts comments
Bio [23] 45730 8/1/1 Fractional area of exposed non polar residue
House [24] 21613 17/1/1 Square footage of the apartments interior living space
Meps19 [2] 15785 138 /3/1 Overall rating of feelings, age, working limitation
NSLM [21] 10391 10/1/1 Synthetic normally distributed random variable
@@ D.2 NSLM dataset details
The 2018 Atlantic Causal Inference Conference workshop on heterogeneous treatment effects [49] introduced the National Study of Learning Mindsets (NSLM) dataset [21]. We refer to [49, Section 2] for its full details. In our experiments, we adopt the approach outlined in [49, 6] to generate synthetic potential outcomes and a synthetic PI variable. The process begins by standardizing the dataset so that all features have a mean of 0 and a standard deviation of 1. The data is then randomly split into two subsets: 80% samples are used for training, while the remaining 20% samples are used for validation. To model the relationship between $X$ and $Y$, we train a neural network with a single hidden layer containing 32 neurons. The learning mechanism of the network is described in Section E.1. We refer to its learned function as $\hat{\mu}_{0}(\cdot)$. Additionally, we employ an XGBoost classifier to estimate the original treatment variable $M$ using the features in $X$. The classifier is configured with a maximum depth of 2 and uses 10 estimators. We then calibrate the estimated propensity score to have the same marginal probability as the original probability and denote the calibrated score $\hat{e}(X_{i})$. Once these models are trained, we generate a new treatment indicator $M_{i}$, a synthetic PI variable $Z_{i}$, and a semi-synthetic outcome variable $Y_{i}$ as detailed next:
@@ D.3 Synthetic dataset details
In this section, we present the synthetic datasets used in this work. Across all datasets, $X,Y,Z$ are generated using the same procedure from [5], where the only difference is the label corruption mechanism. We first describe the generation process for $X,Y,Z$ and then detail the corruption mechanisms.
@@
$\begin{split}\beta_{1}&\sim\text{Uni}(0,1)^{1}\\ \beta_{1}&=\beta_{1}/||\beta_{1}||_{1}\\ \beta_{2}&\sim\text{Uni}(0,1)^{p}\\ \beta_{2}&=\beta_{2}/||\beta_{2}||_{1}\\ {Z^{\prime}}_{i}&=\beta_{2}Z_{i}\\ U_{i}&=\mathbbm{1}_{{Z^{\prime}}_{i}<-3}+2*\mathbbm{1}_{-3\leq{Z^{\prime}}_{i}\leq 1}+8*\mathbbm{1}_{{Z^{\prime}}_{i}>1}\\ E_{i}&\sim\mathcal{N}(0,1)\\ \end{split}$ (109)
@@ E.1 General setup
Across all experiments, the data is divided into a training set (50%), a calibration set (20%), a validation set (10%) used for early stopping, and a test set (20%) for performance evaluation. For the UI method, we further split the original calibration set equally into a reference set (50%) and a calibration set (50%). Next, we normalize the feature vectors and response variables so that they have a zero mean and unit variance. In experiments with missing variables, we impute them using a linear model that is fitted on the variables that are always observed, out of $X,Y,Z$. This linear model is trained using samples from both the training and validation sets. For datasets that are not originally corrupted, the corruption probability is defined as follows. First, for the MEPS19 dataset, we fit a random forest model on the entire dataset to predict the 70% and 30% quantiles of $Y$ given $Z$, and we use their difference as the initial value. For the other datasets, we take $Z$ as the initial value; if $Z$ is multi-dimensional, we multiply it by a random vector to convert it into a scalar. We start by subtracting the minimum value between the 5% quantile and 0, then divide by the 95% quantile of these values, and multiply by 2.5. Negative values are then set to zero. Next, we raise $e$ to the power of the negative of these values and subtract the result from 1. Finally, we zero out any negative values and those below the 77% quantile. To obtain the corruption probabilities, these values are normalized to lie between 0.2 and 0.9 and then raised to a power that results in a 20% marginal corruption probability. Thus, by definition, the average corruption probability is 20%. In every experiment, we train a base learning model and then wrap it with a calibration scheme. The learning model is designed to estimate the 5% and 95% conditional quantiles of $Y\mid X$. In Table 3, we summarize the models employed for each dataset across both tasks. For neural network models, we use an Adam optimizer [50] with a learning rate of 1e-4 and a batch size of 128. The network architecture contains hidden layers with sizes 32, 64, 64, 32, with a dropout rate of 0.1, and uses leaky ReLU as the activation function. We used 100 estimators for both xgboost and random forest models. The networks are trained for 1000 epochs; however, training stops early if the validation loss does not improve for 200 epochs, at which point the model with the lowest validation loss is selected. We used the scikit-learn package [51] to construct random forest models and the xgboost package [51]. The neural networks are implemented using the PyTorch package [52]. The hyperparameters we employed are the default ones unless stated otherwise. For PCP, we set the parameter $\beta$ to $\beta=0.005$. In all experiments in which we employed UI, unless specified otherwise, we used a Full+Linear model for the label regression model $\hat{g}(x,z)$, in which a linear model is given both $Z$ and the output of a neural network model trained using $X,Z$. Moreover, the conditional errors were sampled using the linear clustering approach described in Section C.3.1 When applying a Kmeans clustering, we use the default number of clusters $k=8$.
@@
Facebook1 [22] Neural network Neural network
Facebook2 [22] Neural network Neural network
Bio [23] Neural network Neural network
House [24] Neural network Neural network
Meps19 [2] Random forest Random forest
NSLM [21] XGBoost XGBoost
Synthetic datasets [5] Neural network Neural network
@@ F.3 Causal inference experiment: NSLM dataset
In this causal inference example, our goal is to estimate the uncertainty of individual treatment effects [53]. We utilize the semi-synthetic National Study of Learning Mindsets (NSLM) dataset [21], which deals with behavioral interventions. Further details about the dataset can be found in [49, Section 2], and Appendix D.2 outlines our adaptation for this dataset. Here, ${X}_{i}$ are the individual’s characteristics, $Z_{i}$ are the privileged information, $M_{i}\in\{0,1\}$ denotes the binary treatment indicator, and $Y_{i}(0),Y_{i}(1)\in\mathbb{R}$ denote the counterfactual outcomes under control and treatment conditions, respectively. In practice, we only observe one of them, $\tilde{Y}_{i}$, which equals to $Y_{i}(0)$ if $M_{i}=0$ and to $Y_{i}(1)$ if $M_{i}=1$. In this task, our goal is to estimate the uncertainty of the unknown response under no treatment $Y_{n+1}\equiv Y_{n+1}(0)$ at a pre-specified level $1-\alpha=90\%$. As explained in [5], estimating uncertainty for $Y_{i}(0)$ is crucial since it can be used to construct a reliable prediction interval for the individual treatment effect (ITE), $Y_{i}(1)-Y_{i}(0)$, which is of great interest in many causal inference applications [54, 55, 56, 57]. Furthermore, [5] emphasize that constructing valid prediction sets for $Y_{n+1}(0)$ is challenging due to the distribution shift between the observed control responses, which are drawn from $P_{Y(0)\mid M=0}$. In contrast, the test control response is drawn from $P_{Y(0)}$. Moreover, [5] highlight the difficulty of constructing valid prediction sets for $Y_{n+1}(0)$, as it requires correcting the distribution shift between the observed control responses, which follow the distribution $P_{Y(0)\mid M=0}$, and the test control responses, which are drawn from $P_{Y(0)}$.
@@ F.4 Imputation and error sampling methods
  • Linear: where $\hat{g}$ is a linear function of both $X$ and $Z$, implemented using Scikit-learn’s LinearRegressor [51];

### v3

@@ A.3 Triply robust calibration
Under the assumptions of Theorem 1 we get $\mathbb{P}(Y^{\text{test}}\in C^{\texttt{PCP}}(X^{\text{test}}))\geq 1-\alpha$, under the assumptions of Theorem 4 we get $\mathbb{P}(Y^{\text{test}}\in C^{\texttt{PCP}}(X^{\text{test}}))\geq 1-\alpha$. By Vovk et al. (2005); Angelopoulos & Bates (2023), CP constructs prediction sets with the nominal coverage level when the scores are exchangeable, i.e., $\mathbb{P}(Y^{\text{test}}\in C^{\texttt{CP}}(X^{\text{test}}))\geq 1-\alpha$. Thus, when at least one of the above assumptions hold, one of the above prediction sets achieves the desired coverage rate. Since TriplyRobust is the union of all three sets, its coverage rate is greater than or equal to the coverage rate of each prediction set. Therefore:
@@ Appendix B Additional related work
Practical examples for tasks with privileged information include crowd‑sourcing settings, such as CIFAR-10H, in which annotator metadata, e.g., response times, confidence scores, and expertise level, may serve as PI. Another example is e‑commerce recommendation, where a user’s click history can be considered as PI, which is a good predictor for actual purchases (Yang et al., 2022). In this example, vendors may not share the user’s click history with the model provider due to privacy issues, which makes the PI unavailable at test time. In medical imaging, a pathologist’s diagnostic report for a biopsy image can act as privileged information that strongly predicts whether the tissue is cancerous or healthy (Lopez-Paz et al., 2015). In this context, the PI may not be available for all patients at test time for various reasons, such as limited resources or prioritization.
Several recent works extend conformal prediction beyond the i.i.d. assumption to handle various forms of distribution shifts, including arbitrary shifts (Barber et al., 2023), online and time‐series settings (Gibbs & Candes, 2021; Gibbs & Candès, 2024; Zaffran et al., 2022; Feldman et al., 2023; Xu & Xie, 2023), missing covariates (Zaffran et al., 2023; 2024), and ambiguous labels (Stutz et al., 2023; Caprio et al., 2024). Related research on missing‐data imputation includes analyses of missingness mechanisms (Rubin, 1976), and multiple imputation frameworks (Rubin, 1996; 2018). Furthermore, the work of Zhang (2016) studies regression imputation approaches and discusses adding residual variance to account for prediction uncertainty.
@@ C.1 Privileged conformal prediction
[…]
@@ C.3 Uncertain imputation
Compute the errors $E_{i}$ of $\hat{g}$ on the reference set $\mathcal{I}_{3}$ according to equation 10.
Generate the imputed labels $\bar{Y}_{i},\forall i\in\mathcal{I}_{2}$, according to equation 11.
[…]
@@ C.3.1 Error sampling techniques
In this section, we present our suggestions for sampling errors conditional on the privileged information for the imputation process of UI in equation 11. Instead of using a strict equality condition of $Z_{i}=z$ in $\mathcal{E}(z):=\{E_{i}:i\in\mathcal{I}_{3},Z_{i}=z\}$, we can relax this requirement by clustering the $Z$ space and comparing the clusters into which each $z$ falls, that is, $\mathcal{E}(z):=\{E_{i}:i\in\mathcal{I}_{3},h(Z_{i})=h(z)\}$, where $h$ is a clustering function. The first clustering method we use is Kmeans, which we fit on the training and validation data. Then, we cluster the PIs of the reference set. When imputing the corrupted labels of the calibration set, we sample an error from the cluster corresponding to the test PI. The second clustering approach we examine is Linear clustering. We first fit on the training and validation data a linear model that takes as an input the PI $Z$ and outputs an estimated label $Y$. We used this model to compute the estimated labels for each point in the reference set and then split them into bins. We note that the alternative approach of learning the error distribution given the PI $Z$ may be applied as well, such as random forest for conditional density estimation (Pospisil & Lee, 2018), or normalizing flows (Rezende & Mohamed, 2015; Papamakarios et al., 2021).
@@ C.4 Method validity conditions summary
Quantile Regression (Koenker & Bassett, 1978; Koenker, 2005; Steinwart & Christmann, 2011; Takeuchi et al., 2006) $\checkmark$ NA NA
@@ D.1 General real dataset details
Dataset # Samples $\bm{X/Z/Y}$ Dimensions $\bm{Z}$ description
facebook1 (facebook, ) 40948 52/1/1 Number of posts comments
facebook2 (facebook, ) 81311 52/1/1 Number of posts comments
Bio (bio, ) 45730 8/1/1 Fractional area of exposed non polar residue
House (house, ) 21613 17/1/1 Square footage of the apartments interior living space
Meps19 (meps_, 19) 15785 138 /3/1 Overall rating of feelings, age, working limitation
NSLM (Yeager et al., 2019) 10391 10/1/1 Synthetic normally distributed random variable
@@ D.2 NSLM dataset details
The 2018 Atlantic Causal Inference Conference workshop on heterogeneous treatment effects (Carvalho et al., 2019) introduced the National Study of Learning Mindsets (NSLM) dataset (Yeager et al., 2019). We refer to Carvalho et al. (2019, Section 2) for its full details. In our experiments, we adopt the approach outlined in Carvalho et al. (2019); Lei & Candès (2021) to generate synthetic potential outcomes and a synthetic PI variable. The process begins by standardizing the dataset so that all features have a mean of 0 and a standard deviation of 1. The data is then randomly split into two subsets: 80% samples are used for training, while the remaining 20% samples are used for validation. To model the relationship between $X$ and $Y$, we train a neural network with a single hidden layer containing 32 neurons. The learning mechanism of the network is described in Section E.1. We refer to its learned function as $\hat{\mu}_{0}(\cdot)$. Additionally, we employ an XGBoost classifier to estimate the original treatment variable $M$ using the features in $X$. The classifier is configured with a maximum depth of 2 and uses 10 estimators. We then calibrate the estimated propensity score to have the same marginal probability as the original probability and denote the calibrated score $\hat{e}(X_{i})$. Once these models are trained, we generate a new treatment indicator $M_{i}$, a synthetic PI variable $Z_{i}$, and a semi-synthetic outcome variable $Y_{i}$ as detailed next:
@@ D.3 Synthetic dataset details
In this section, we present the synthetic datasets used in this work. Across all datasets, $X,Y,Z$ are generated using the same procedure from Feldman & Romano (2024), where the only difference is the label corruption mechanism. We first describe the generation process for $X,Y,Z$ and then detail the corruption mechanisms.
@@
[…]
@@ E.1 General setup
Across all experiments, the data is divided into a training set (50%), a calibration set (20%), a validation set (10%) used for early stopping, and a test set (20%) for performance evaluation. For the UI method, we further split the original calibration set equally into a reference set (50%) and a calibration set (50%). Next, we normalize the feature vectors and response variables so that they have a zero mean and unit variance. In experiments with missing variables, we impute them using a linear model that is fitted on the variables that are always observed, out of $X,Y,Z$. This linear model is trained using samples from both the training and validation sets. For datasets that are not originally corrupted, the corruption probability is defined as follows. First, for the MEPS19 dataset, we fit a random forest model on the entire dataset to predict the 70% and 30% quantiles of $Y$ given $Z$, and we use their difference as the initial value. For the other datasets, we take $Z$ as the initial value; if $Z$ is multi-dimensional, we multiply it by a random vector to convert it into a scalar. We start by subtracting the minimum value between the 5% quantile and 0, then divide by the 95% quantile of these values, and multiply by 2.5. Negative values are then set to zero. Next, we raise $e$ to the power of the negative of these values and subtract the result from 1. Finally, we zero out any negative values and those below the 77% quantile. To obtain the corruption probabilities, these values are normalized to lie between 0.2 and 0.9 and then raised to a power that results in a 20% marginal corruption probability. Thus, by definition, the average corruption probability is 20%. In every experiment, we train a base learning model and then wrap it with a calibration scheme. The learning model is designed to estimate the 5% and 95% conditional quantiles of $Y\mid X$. In Table 3, we summarize the models employed for each dataset across both tasks. For neural network models, we use an Adam optimizer (Kingma & Ba, 2015) with a learning rate of 1e-4 and a batch size of 128. The network architecture contains hidden layers with sizes 32, 64, 64, 32, with a dropout rate of 0.1, and uses leaky ReLU as the activation function. We used 100 estimators for both xgboost and random forest models. The networks are trained for 1000 epochs; however, training stops early if the validation loss does not improve for 200 epochs, at which point the model with the lowest validation loss is selected. We used the scikit-learn package (Pedregosa et al., 2011) to construct random forest models and the xgboost package (Pedregosa et al., 2011). The neural networks are implemented using the PyTorch package (Paszke et al., 2019). The hyperparameters we employed are the default ones unless stated otherwise. For PCP, we set the parameter $\beta$ to $\beta=0.005$. In all experiments in which we employed UI, unless specified otherwise, we used a Full+Linear model for the label regression model $\hat{g}(x,z)$, in which a linear model is given both $Z$ and the output of a neural network model trained using $X,Z$. Moreover, the conditional errors were sampled using the linear clustering approach described in Section C.3.1 When applying a Kmeans clustering, we use the default number of clusters $k=8$.
@@
Facebook1 (facebook, ) Neural network Neural network
Facebook2 (facebook, ) Neural network Neural network
Bio (bio, ) Neural network Neural network
House (house, ) Neural network Neural network
Meps19 (meps_, 19) Random forest Random forest
NSLM (Yeager et al., 2019) XGBoost XGBoost
Synthetic datasets (Feldman & Romano, 2024) Neural network Neural network
@@ F.3 Causal inference experiment: NSLM dataset
In this causal inference example, our goal is to estimate the uncertainty of individual treatment effects (Hernán & Robins, 2010). We utilize the semi-synthetic National Study of Learning Mindsets (NSLM) dataset (Yeager et al., 2019), which deals with behavioral interventions. Further details about the dataset can be found in Carvalho et al. (2019, Section 2), and Appendix D.2 outlines our adaptation for this dataset. Here, ${X}_{i}$ are the individual’s characteristics, $Z_{i}$ are the privileged information, $M_{i}\in\{0,1\}$ denotes the binary treatment indicator, and $Y_{i}(0),Y_{i}(1)\in\mathbb{R}$ denote the counterfactual outcomes under control and treatment conditions, respectively. In practice, we only observe one of them, $\tilde{Y}_{i}$, which equals to $Y_{i}(0)$ if $M_{i}=0$ and to $Y_{i}(1)$ if $M_{i}=1$. In this task, our goal is to estimate the uncertainty of the unknown response under no treatment $Y_{n+1}\equiv Y_{n+1}(0)$ at a pre-specified level $1-\alpha=90\%$. As explained in Feldman & Romano (2024), estimating uncertainty for $Y_{i}(0)$ is crucial since it can be used to construct a reliable prediction interval for the individual treatment effect (ITE), $Y_{i}(1)-Y_{i}(0)$, which is of great interest in many causal inference applications (Brand & Xie, 2010; Morgan, 2001; Xie et al., 2012; Florens et al., 2008). Furthermore, Feldman & Romano (2024) emphasizes that constructing valid prediction sets for $Y_{n+1}(0)$ is challenging due to the distribution shift between the observed control responses, which are drawn from $P_{Y(0)\mid M=0}$. In contrast, the test control response is drawn from $P_{Y(0)}$. Moreover, Feldman & Romano (2024) highlight the difficulty of constructing valid prediction sets for $Y_{n+1}(0)$, as it requires correcting the distribution shift between the observed control responses, which follow the distribution $P_{Y(0)\mid M=0}$, and the test control responses, which are drawn from $P_{Y(0)}$.
@@ F.4 Imputation and error sampling methods
  • Linear: where $\hat{g}$ is a linear function of both $X$ and $Z$, implemented using Scikit-learn’s LinearRegressor (Pedregosa et al., 2011);
@@
F.5 Empirically evaluating the assumptions of Theorem 4
In this section, we assess whether the assumptions of Theorem 4 are satisfied in practice, by conducting two experiments following the protocol in Section 4.3.
In the first experiment, we evaluate the dependence between residual errors $Y^{\text{test}}-\hat{g}(X^{\text{test}},Z^{\text{test}})$ and (i) the predictions $\hat{g}(X^{\text{test}},Z^{\text{test}})$, (ii) the lower bounds $C^{0}(X^{\text{test}})$, and (iii) the upper bounds $C^{1}(X^{\text{test}})$ of the intervals produced by UI. We assess the dependence by computing the partial correlation (PC) conditional $Z^{\text{test}}$ between the terms. We emphasize that residual errors were used in place of $R^{\text{test}}$ because the latter is unavailable in practice. Likewise, since the true function $g^{*}$ is unknown, we replaced it with its estimator $\hat{g}$. These replacements serve as fair alternatives for the unknown variables. For completeness, we also report the coverage rate attained by UI and the prediction error (MSE) of $\hat{g}$. All metrics were computed on the test set, and the results are averaged over 10 random data splits. The empirical results, summarized in Table 4, reveal that there is a non-negligible partial correlation between the interval endpoints and the residual errors. This suggests that the independence requirements of Theorem 4 are not exactly satisfied. Nevertheless, the intervals of UI still achieve the nominal coverage rate. This observation reveals the robustness of the UI procedure: even when the theoretical conditions are only approximately satisfied, it still constructs valid intervals.
Table 4: Metrics assessing the assumptions of Theorem 4.
Dataset Coverage $\hat{g}$ MSE $\hat{g}$ PC $C^{0}$ PC $C^{1}$ PC
Bio $90.72\pm 0.49$ $56\pm 0.00 0.0$$2\pm 0.00 0.0$$3\pm 0.00 0.0$$02\pm 0.00$
Facebook1 $90.96\pm 0.36$ $1.84\pm 0.21$ $07\pm 0.05 0.0$$2\pm 0.04 0.0$$03\pm 0.05$
House $91.23\pm 0.36$ $0.71\pm 0.03$ $0.15\pm 0.01$ $0.18\pm 0.01$ $0.19\pm 0.01$
In the second experiment, we analyze the setup where the PI is a weak predictor of $Y$. We follow the same experimental protocol in Section 4.3 and simulate this setup by adding random Gaussian noise to $Z$, with varying standard deviations. Figure 15 summarizes the performance of UI with these weak PIs, indicating that as the magnitude of the noise increases, the coverage rate of UI increases. That is, as the PI becomes a weak predictor of the labels, the uncertainty of the imputed labels increases, which, in turn, widens the resulting prediction sets. This experiment suggests that UI can still obtain a valid coverage rate even when the PI is a weak predictor of the labels.
Figure 15: Weak PI experiment. The coverage rate as a function of the noise of the PI. The coverage rate is evaluated over 4 random data splits.
F.6 The performance of TriplyRobust
We employ TriplyRobust, which combines CP, PCP that uses either the estimated corruption probabilities or the true ones, and UI on the datasets from Section 4.3. We present the coverage rate and interval length in Figure 16. This figure shows that TriplyRobust employed with PCP that uses estimated weights constructs wider intervals, which is in line with the results in Section 4.3, in which this version of PCP constructs wider intervals, as its approximations are not sufficiently accurate. Nevertheless, this figure reveals that combining the three approaches does not significantly harm the statistical efficiency of the predicted intervals, especially when using oracle or sufficiently accurate weights.
Figure 16: Missing response experiment. The coverage rate and average interval length obtained by TriplyRobust employed with PCP that uses either the estimated corruption probabilities or the true ones. Performance metrics are evaluated over 30 random data splits.